# Evaluation of the absolute single-photon detection efficiency of HRPPD

### v1

@@ Abstract
Pixelated High Rate Picosecond Photon Detectors (HRPPDs) by Incom Inc. are promising photosensors for use in Ring Imaging CHerenkov (RICH) detectors, where a high gain, sub-mm position resolution and sub-100ps timing resolution are required in a single photon mode. Quantum Efficiency (QE) has been measured for the first batch of EIC HRPPDs both by Incom and EIC research groups at Jefferson Lab and Brookhaven Lab, with peak values at $\sim$365 nm typically exceeding 30% ([1]). In this study, we present a first direct measurement of HRPPD Photon Detection Efficiency (PDE) being equal to 19.3 $\pm$ 0.2 (stat) $\pm$ 0.8 (sys) % at 398.6 nm, in a photoelectron pulse counting mode using a picosecond diode laser. HRPPD QE at the same spot and at the same wavelength was evaluated to be 24.3 $\pm$ 0.1 (stat) $\pm$ 0.5 (sys) %, leading to a Collection Efficiency (CE) estimate of $\sim$79%, which is consistent with the expectations.
@@ 1 Methodology
$P(\text{no signal})=\sum_{n=0}^{\infty}P(n)(1-p)^{n},$ (1)
@@
The mean number of photons in a single laser pulse ($\lambda$) incident on HRPPD can be derived from the following expression: $\lambda=\frac{I\cdot R}{e\cdot QE\cdot f\cdot A}$, where $I$ is the current as measured by the reference Hamamatsu photodiode (see Figure 1), QE is its quantum efficiency at the laser wavelength, $R$ is a ratio of the transmitted and the reflected light after the beam splitter at this wavelength, $e$ is the elementary charge, $f$ is a laser repetition rate, and $A$ is an attenuation factor of the ND filter.
@@
It is worth noting that the photon detection efficiency (PDE) is typically determined using a different method ([2]). In that approach, the pulse amplitude at the anode is measured and then divided by the gain, obtained from a spectral fit, to estimate the number of photoelectrons generated after the photocathode. In contrast, we count signal pulses in a binary manner, thereby eliminating the dependence on gain.
@@ 2.1 HRPPD and readout backplane
We used HRPPD #16 (serial number assigned by Incom Inc.) for these measurements. It was mounted on a side wall of the dark box, as shown in Figure 3. The HRPPD was equipped with a Q02c passive readout backplane and a pair of 32-channel M02c MCX adapters, which allowed us to fully instrument a single 8x8 anode pad spot.
@@
Figure 3: HRPPD #16 mounted on a dark box side wall. One can also see the interface board with a pair of plugin MCX adapters.
@@ 2.2 Dark box equipment
A UV-resilient UM22-600 multi-mode fiber with a core diameter of 600 $\mu$m by Thorlabs was used between the dark box patch panel and the optical head for both PDE and QE, as well as for the calibration measurements. The light was focused by a LA1422 N-BK7 plano-convex lens with a focal distance of 40 mm, with its intensity and angular spread limited by a SM1D12C diaphragm with a chosen $\sim$3.5 mm diameter setting, and split into two paths by a CM1-BP145B5 pellicle beam splitter, all by Thorlabs.
@@
None of the elements shown in Figure 4 was either changed or moved between PDE and QE measurements. In particular, we used the same fiber, diaphragm diameter setting, and XYZ-coordinates of the optical head. For calibration, either an ND filter or a Newport photodiode was mounted downstream of the beam splitter as explained in the subsequent sections.
@@ 2.3 High Voltage
For the PDE measurements, the HRPPD was biased using a stackable CAEN A1515BV high voltage power supply in a SY5527LC mainframe. We used the following settings: MCP bias 650 V, photocathode and transfer voltages: 200 V (gain around $4\times 10^{6}$ for a particular illuminated pixel, see also a blue dashed line in Figure 10).
@@ 2.6 Spectrometer
Figure 5: Left: PiLas laser wavelength spectrum with a peak value of 398.4 nm and FWHM 1.9 nm as measured by the Ocean Optics spectrometer. Right: Newport monochromator wavelength spectrum with a FWHM 3.0 nm and a peak value matching the laser one within 0.1 nm as measured by the same spectrometer when using a monochromator wavelength setting of 395.7 nm .
@@
As follows from Figure 5 (left), a measured PiLas laser wavelength of 398.4 nm matched very well 398.6 nm provided in the factory report. However, the monochromator scale was found to be off by almost 3 nm. In the complementary QE measurements described in Section 5, we therefore used a slightly different setting of 395.7 nm in the monochromator steering script, in order to perfectly match the wavelength spectra peak values between the PiLas laser and the monochromator, as seen in Figure 5.
@@ 2.7 Picoammeters
A Keithley 6485 unit was used to measure the photocurrent of the Hamamatsu reference photodiode. A Keithley 6487 was used for the NIST traceable Newport photodiode as well as the HRPPD. For correlated measurements between a pair of picoammeters (see Section 3 below) their measurement ranges have always been fixed to a value matching the largest measured current, and the AUTO range setting feature turned off.
@@ 2.8 Beam spot imaging
We used an AmScope MD310C-BS bare sensor CMOS camera with a 5.12 mm x 3.84 mm sensor (2.5 $\mu$m pixels), installed at a location of the HRPPD window to air boundary, to verify a focused beam spot size. Obtained images for both the laser and the monochromator light sources are shown in Figure 6. As seen from this figure, beam diameter in both cases was about 1.0 mm (mostly defined by the fiber diameter of 600 $\mu$m), though the monochromator image looks like a ”more smooth” disc. A beam spot of this size was indeed well within the 8 mm x 8 mm active area of the Hamamatsu reference photodiode.
Figure 6: Left: PiLas laser beam spot image at the HRPPD window to air boundary location. Right: Newport monochromator beam image in the same optical configuration (fiber type, lens location, diaphragm opening).
@@ 2.9 DAQ, readout electronics, steering software
Software used to steer all of the setup components, as well as the RCDAQ data acquisition system was installed on a Debian 12 desktop PC. Python scripts to operate the monochromator, picoammeters, and translation stages, were written from scratch. Trigger logic was assembled in a NIM crate, while a pair of CAEN V1742 digitizers in a VME crate was operated via respective CAENET optical links by RCDAQ via a PCIe A3818 interface card.
@@ 3 Reference photodiode and ND filter calibration
The reference Hamamatsu photodiode does not come with a reliable factory calibration. Besides this, as it was mentioned earlier, neither the beam splitter ratio nor the ND filter optical density at a laser wavelength is known with a sufficient accuracy. In order to relate a reference Hamamatsu photodiode current to a NIST traceable Newport photodiode one, in a configuration with a beam splitter and (optionally) an ND filter installed downstream of it, we performed a series of measurements in configurations shown schematically in Figure 7. We considered three cases: (1) no filter installed between the beam splitter and the Newport photodiode, (2) a nominally OD3 reflective filter installed which came together with the Newport photodiode (see Section 2.2), (3) a nominally OD4 absorptive Thorlabs filter we were eventually using to perform the PDE measurements.
Figure 7: Configurations used to cross-calibrate Newport and Hamamatsu photodiodes, as well as our OD4 neutral density filter.
Results are presented in Figure 8. The PiLas Laser was set at a 0% tune (maximum pulse intensity). Average integrated light intensity was varied by changing a laser repetition rate between 30 kHz and 40 MHz in roughly half a decade steps (10 points on a ”no filter” curve in Figure 8). Ten current measurements per point were taken, and their average and standard deviation are used in three separate $f(x)=ax+b$ straight line fits. In all three cases, we observed a perfect linearity within the measurement errors ($\chi^{2}$ of 5.4, 5.2, and 0.02 per 8, 6, and 2 degrees of freedom, respectively), and $b$-terms consistent with 0.
Figure 8: Currents measured by the Newport and Hamamatsu photodiodes at various PiLas laser intensities, for the configurations shown in Figure 7.
We used this evidence as an indirect proof that (1) Hamamatsu photodiode has a linear response over four decades in light intensity, (2) one can safely use a slope of the ”no filter” straight line in Figure 8 to calibrate the Hamamatsu reference photodiode against the Newport one for QE measurements, (3) in a similar way, one can use a ratio of the ”Our OD4 filter” and ”no filter” line slopes in Figure 8 to cross-calibrate the Hamamatsu photodiode in a low light intensity configuration required to measure the HRPPD PDE in a photoelectron counting mode. Numerically, the slope of the ”no filter” straight line was found equal to 0.8409 $\pm$ 0.0003, and the ratio of the ”no filter” and ”Our OD4 filter” slopes (effectively the optical density of our OD4 filter at 398.6 nm) was equal to 147420 $\pm$ 90.
It is worth mentioning that the ”Newport OD3 filter” data were taken with a neutral density filter, for which a separate factory calibration was provided (see Section 2.2). Our measured slope ratio of the ”Newport OD3 filter” and ”no filter” curves came out equal to 416.9 $\pm$ 0.2, consistent with these factory calibrations at 398.6 nm within the quoted errors.
@@ 4 HRPPD PDE measurement
The measurement was performed in a configuration shown schematically in Figure 1, with our nominally OD4 neutral density filter installed between the beam splitter and the HRPPD window. The PiLas laser was run with a repetition rate of 1 kHz at a set of tunes listed in Table 1. RCDAQ DAQ was used to take the waveform data of all the 8x8 pads around the illuminated spot. V1742 digitizers were running at a 5 GS/s sampling rate. Trigger for the DAQ was provided by the PiLas laser itself. A panel of a typical ”event” is shown in Figure 9, where a horizontal range is 40 ns and a vertical range is 50 mV in each of the waveforms of the 3x3 pads surrounding the illuminated one.
@@
Despite the fact that a majority of recorded events looked like in Figure 9 (only the illuminated pad had a measurable signal), we also considered various summation schemes for a 3x3 pad area around the illuminated spot. At the end, we decided to use waveforms of the illuminated pad and four neighboring pads for the collected charge measurements. No peak search algorithm was used. HRPPD signal was taken as a sum of 16 waveform samples (a 3.2 ns wide window) at a fixed timing offset with respect to a digitized trigger pulse. The baseline was estimated on an event-per-event basis in a same width (3.2 ns) window just in front of the location of the expected HRPPD peak, in the same waveform. Collected charge was calculated as a difference between the ”peak” value (a sum over 16 samples as explained earlier, whether there was a measurable peak observed or not) and a baseline. Charge distribution, normalized to a detected electron count (which in a pure single photon mode would effectively have a meaning of the HRPPD gain) is shown in Figure 10, together with a Gaussian fit to a pedestal peak (events where HRPPD did not produce a measurable signal) and a two-Gaussian fit to the actual signal distribution, assuming that three-photoelectron event probability was small and therefore irrelevant.
Figure 10: Charge collected by the illuminated HRPPD pad. Red dashed line indicates the fit to the pedestal, blue dashed line the one-photon peak. Blue solid line is the sum of the fits to one-photon peak and two-photon peak.
One can see that the HRPPD was operated at a gain of about $4\times 10^{6}$. The pedestal peak had $\sigma\sim 0.09\times 10^{6}$. A fraction of events with a measurable HRPPD signal required in formula 3 was determined as a ratio of events with a collected charge above $3\sigma$ with respect to the pedestal mean, and the total number of triggers.
@@
We repeated the same measurements at a set of different laser tunes corresponding to a substantially different average fraction of non-empty events, and obtained consistent PDE values, all listed in Table 1. A statistical average of 19.3 $\pm$ 0.2 ($\chi^{2}\sim 2.2$ for 2 degrees of freedom) is taken as a final result.
tune $\lambda$ events above threshold PDE
36.0% 4.978 $\pm$ 0.010 61.4 $\pm$ 0.4% 18.9 $\pm$ 0.2%
46.0% 3.713 $\pm$ 0.007 51.7 $\pm$ 0.5% 19.4 $\pm$ 0.3%
52.4% 2.573 $\pm$ 0.013 39.8 $\pm$ 0.4% 19.5 $\pm$ 0.3%
Table 1: Results of three sets of data using different laser tunes, listed in the first column. The second column ($\lambda$) is the mean number of photons impinging on the HRPPD photocathode per laser pulse. Only statistical errors are quoted.

### v2

@@ Abstract
Pixelated High Rate Picosecond Photon Detectors (HRPPDs) by Incom Inc. are promising photosensors for use in Ring Imaging CHerenkov (RICH) detectors, where a high gain, sub-mm position resolution and sub-100ps timing resolution are required in a single photon mode. Quantum Efficiency (QE) has been measured for the first batch of EIC HRPPDs both by Incom and EIC research groups at Jefferson Lab and Brookhaven Lab, with peak values at $\sim$365 nm typically exceeding 30%. In this study, we present a first direct measurement of HRPPD Photon Detection Efficiency (PDE) being equal to 17.13 $\pm$ 0.11 (stat) $\pm$ 0.29 (sys) % at 398.6 nm, in a photoelectron pulse counting mode using a picosecond diode laser. HRPPD QE at the same spot and at the same wavelength was evaluated to be 24.37 $\pm$ 0.03 (stat) $\pm$ 0.32 (sys) %, leading to a Collection Efficiency (CE) estimate of 70.3 $\pm$ 1.6%, which is consistent with the expectations.
@@
1 Introduction
Pixelated High Rate Picosecond Photon Detectors (HRPPDs), a new type of microchannel plate photomultiplier tube, thanks to their high gain, excellent timing resolution, and sub-mm position resolution, serve as promising candidates for use in Ring Imaging CHerenkov (RICH) detectors at the future EIC [1] and upgrades for LHCb [2] and Belle II [3].
Recently, Quantum Efficiency (QE) of the first batch of EIC HRPPDs produced by Incom Inc. [4] has been measured both by Incom and EIC research groups at Jefferson Lab and Brookhaven Lab, with peak values at $\sim$365 nm typically exceeding 30% [5]. In addition to QE, the photon detection efficiency (PDE) provides a more direct and practical measure of the device’s feasibility for use in RICH detectors. In this paper, we present the results of the first measurement of the absolute photon detection efficiency of HRPPDs using a novel approach.
@@ 2 Methodology
$P(\text{no signal})=\sum_{n=0}^{\infty}P(n)(1-p)^{n}$ (1)
@@
The mean number of photons in a single laser pulse ($\lambda$) incident on HRPPD can be derived from the following expression:
$\lambda=\frac{I\cdot R}{e\cdot QE\cdot f\cdot A}$ (4)
where $I$ is the current as measured by the reference Hamamatsu photodiode (see Figure 1), QE is its quantum efficiency at the laser wavelength, $R$ is a ratio of the transmitted and the reflected light after the beam splitter at this wavelength, $e$ is the elementary charge, $f$ is a laser repetition rate, and $A$ is an attenuation factor of the ND filter.
@@
It is worth noting that the photon detection efficiency (PDE) is typically determined using a different method [6]. In that approach, the pulse amplitude at the anode is measured and then divided by the gain, obtained from a spectral fit, to estimate the number of photoelectrons generated after the photocathode. In contrast, we count signal pulses in a binary manner, thereby eliminating the dependence on gain.
@@ 3.1 HRPPD and readout backplane
We used HRPPD #27 (serial number assigned by Incom Inc.) for these measurements. It was mounted on a side wall of the dark box, as shown in Figure 3. The HRPPD was equipped with a Q02c passive readout backplane and a pair of 32-channel M02c MCX adapters, which allowed us to fully instrument a single 8x8 anode pad spot.
@@
Figure 3: HRPPD #27 mounted on a dark box side wall. One can also see the interface board with a pair of plugin MCX adapters.
@@ 3.2 Dark box equipment
A UV-resilient UM22-600 multi-mode fiber with a core diameter of 600 $\mu$m by Thorlabs was used between the dark box patch panel and the optical head for both PDE and QE, as well as for the calibration measurements. The light was focused by a LA1422 N-BK7 plano-convex lens with a focal distance of 50 mm, with its intensity and angular spread limited by a Thorlabs SM1D12C diaphragm with a chosen $\sim$6 mm diameter setting, and split into two paths by a 1 mm thick sapphire plate by MSE Supplies used as an asymmetric (about 83:17) beam splitter with a smooth wavelength dependency.
@@
All elements shown in Figure 4 remain untouched and unmoved, except the ND filter between PDE and QE measurements. In particular, we used the same fiber, diaphragm diameter setting, and XYZ-coordinates of the optical head. For calibration, either an ND filter or a Newport photodiode was mounted downstream of the beam splitter as explained in the subsequent sections.
@@ 3.3 High Voltage
For the PDE measurements, the HRPPD was biased using a stackable CAEN A1515BV high voltage power supply in a SY5527LC mainframe. We used the following settings: MCP bias 600 V, photocathode and transfer voltages: 200 V (gain around $5\times 10^{6}$ for a particular illuminated pixel, see also a blue dashed line in Figure 10).
@@ 3.6 Spectrometer
Figure 5: Top: PiLas laser wavelength spectrum with a peak value of 398.4 nm and FWHM 1.9 nm as measured by the Ocean Optics spectrometer. Bottom: Newport monochromator wavelength spectrum with a FWHM 3.0 nm and a peak value matching the laser one within 0.1 nm as measured by the same spectrometer when using a monochromator wavelength setting of 395.7 nm .
@@
As follows from Figure 5 (left), a measured PiLas laser wavelength of 398.4 nm matched very well 398.6 nm provided in the factory report. However, the monochromator scale was found to be off by almost 3 nm. In the complementary QE measurements described in Section 6, we therefore used a slightly different setting of 395.7 nm in the monochromator steering script, in order to match the wavelength spectra peak values between the PiLas laser and the monochromator, as seen in Figure 5.
@@ 3.7 Picoammeters
A Keithley 6485 unit was used to measure the photocurrent of the Hamamatsu reference photodiode. A Keithley 6487 was used for the NIST traceable Newport photodiode as well as the HRPPD. Data reading is configured in “SLOW” mode, which corresponds to an integration time of one power line cycle ($\sim$ 16.7 ms).
@@ 3.8 Beam spot imaging
We used an AmScope MD310C-BS bare sensor CMOS camera with a 5.12 mm x 3.84 mm sensor (2.5 $\mu$m pixels), installed at a location of the HRPPD window to air boundary, to verify a focused beam spot size. Obtained images for both the laser and the monochromator light sources are shown in Figure 6. As seen from this figure, beam diameter in both cases was about 1.0 mm (mostly defined by the fiber diameter of 600 $\mu$m), though the monochromator image looks like a “smoother” disc. A beam spot of this size was indeed well within the 8 mm x 8 mm active area of the Hamamatsu reference photodiode.
Figure 6: Left: PiLas laser beam spot image at the HRPPD window to air boundary location. Right: Newport monochromator beam image in the same optical configuration (fiber type, lens location, diaphragm opening). Sensor dimension: 5.12 mm x 3.84 mm.
@@ 3.9 DAQ, readout electronics, steering software
Software used to steer all of the setup components, as well as the RCDAQ data acquisition system [7] was installed on a Debian 12 desktop PC. Python scripts to operate the monochromator, picoammeters, and translation stages, were written using PySerial library. Trigger logic was assembled in a NIM crate, while a pair of CAEN V1742 digitizers in a VME crate was operated via respective CAENET optical links by RCDAQ via a PCIe A3818 interface card.
@@ 4 Reference photodiode and ND filter calibration
The reference Hamamatsu photodiode does not come with a reliable factory calibration. Besides this, as it was mentioned earlier, neither the beam splitter ratio nor the ND filter optical density at a laser wavelength is known with a sufficient accuracy. In order to relate a reference Hamamatsu photodiode current to a NIST traceable Newport photodiode one, in a configuration with a beam splitter and (optionally) an ND filter installed downstream of it, we performed a series of measurements in configurations shown schematically in Figure 7. We considered three cases: (1) no filter installed between the beam splitter and the Newport photodiode, (2) a nominally OD3 reflective filter installed which came together with the Newport photodiode (see Section 3.2), (3) a nominally OD5 absorptive Thorlabs filter we were eventually using to perform the PDE measurements.
Figure 7: Configurations used to cross-calibrate Newport and Hamamatsu photodiodes, as well as our OD5 neutral density filter.
Results are presented in Figure 8. The PiLas Laser was set at a 0% tune (maximum pulse intensity). Average integrated light intensity was varied by changing a laser repetition rate decreasing from 40 MHz in roughly half a decade steps (11 points on a “no filter” curve in Figure 8). Ten current measurements per point were taken, and their average and standard deviation are used in three separate $f(x)=ax+b$ straight line fits. In all three cases, we observed a perfect linearity within the measurement errors ($\chi^{2}$ of 3.12, 0.54, and 0.39 per 9, 6, and 8 degrees of freedom, respectively), and $b$-terms consistent with 0. (Note that the small values of $\chi^{2}$ are attributed to the overestimated uncertainties.)
Figure 8: Currents measured by the Newport and Hamamatsu photodiodes at various PiLas laser light intensities, for the configurations shown in Figure 7.
Following observations can be concluded: (1) Hamamatsu photodiode has a linear response over four decades in light intensity, (2) one can use a slope of the “no filter” straight line in Figure 8 to calibrate the Hamamatsu reference photodiode against the Newport one for QE measurements, (3) in a similar way, one can use a ratio of the “OD5 filter” and “no filter” line slopes in Figure 8 to cross-calibrate the Hamamatsu photodiode in a low light intensity configuration required to measure the HRPPD PDE in a photoelectron counting mode. Numerically, the slope of the “no filter” straight line was found equal to 5.0416 $\pm$ 0.0011, and the ratio of the “no filter” and “OD5 filter” slopes (effectively the optical density of the OD5 filter at 398.6 nm) was equal to (1556 $\pm$ 5)*1E3.
It is worth mentioning that the “Newport OD3 filter” data were taken with a neutral density filter, for which a separate factory calibration was provided (see Section 3.2). Our measured slope ratio of the “Newport OD3 filter” and “no filter” curves leads to an attenuation factor of 411.0 $\pm$ 0.2, consistent with these factory calibrations at 398.6 nm within the quoted errors.
@@ 5 HRPPD PDE measurement
The measurement was performed in a configuration shown schematically in Figure 1, with the nominally OD5 neutral density filter installed between the beam splitter and the HRPPD window. The PiLas laser was run with a repetition rate of 1 kHz at a set of tunes listed in Table 1. RCDAQ DAQ was used to take the waveform data of all the 8x8 pads around the illuminated spot. V1742 digitizers were running at a 5 GS/s sampling rate. Trigger for the DAQ was provided by the PiLas laser itself. A panel of a typical “event” is shown in Figure 9, where a horizontal range is 40 ns and a vertical range is 50 mV in each of the waveforms of the 3x3 pads surrounding the illuminated one.
@@
Despite the fact that a majority of recorded events looked like in Figure 9 (only the illuminated pad had a measurable signal), various summation schemes for a 3x3 pad area around the illuminated spot have been tested. Ultimately, we decided to use waveforms of the illuminated pad and four neighboring pads for the collected charge measurements, which contain most part of the signal charge and provide a relatively narrow pedestal peak. No peak search algorithm was used. HRPPD signal was taken as a sum of 16 waveform samples (a 3.2 ns wide window) at a fixed timing offset with respect to a digitized trigger pulse. The baseline was estimated on an event-per-event basis in a same width (3.2 ns) window just in front of the location of the expected HRPPD peak, in the same waveform. Collected charge was calculated as a difference between the “peak” value (a sum over 16 samples as explained earlier, whether there was a measurable peak observed or not) and a baseline. Charge distribution, normalized to a detected electron count (which in a pure single photon mode would effectively have a meaning of the HRPPD gain) is shown in Figure 10, together with a Gaussian fit to a pedestal peak (events where HRPPD did not produce a measurable signal) and a two-Gaussian fit to the actual signal distribution, assuming that three-photoelectron event probability was small and therefore irrelevant.
Figure 10: Charge collected by the illuminated HRPPD pad. Red dashed line indicates a fit to the pedestal, blue dashed line the one-photon peak. Blue solid line is the sum of the fits to one-photon peak and two-photon peak.
One can see that the HRPPD was operated at a gain of about $5\times 10^{6}$. The pedestal peak had $\sigma\sim 0.09\times 10^{6}$. A fraction of events with a measurable HRPPD signal required in formula 3 was determined as a ratio of events with a collected charge above $3\sigma$ with respect to the pedestal mean, and the total number of triggers.
@@
We repeated the same measurements at a set of different laser tunes corresponding to a substantially different average fraction of non-empty events, and obtained consistent PDE values, all listed in Table 1. A statistical average of 17.13 $\pm$ 0.11 ($\chi^{2}/$NDF = 0.87/4, p-value = 0.9283) is taken as a final result.
tune $\lambda$ signal / total PDE
0% 3.9617$\pm$0.0074 49.09$\pm$0.35 % 17.04$\pm$0.22 %
10% 3.3681$\pm$0.0074 43.99$\pm$0.35 % 17.21$\pm$0.23 %
15% 2.9682$\pm$0.0079 40.13$\pm$0.35 % 17.28$\pm$0.24 %
25% 2.3003$\pm$0.0050 32.42$\pm$0.33 % 17.03$\pm$0.23 %
40% 1.6776$\pm$0.0057 24.95$\pm$0.31 % 17.11$\pm$0.26 %
Table 1: Results of five sets of data using different laser tunes, listed in the first column. The second column ($\lambda$) is the mean number of photons impinging on the HRPPD photocathode per laser pulse. Only statistical errors are quoted.